# Type Ia supernova growth-rate measurement with LSST simulations: intrinsic scatter systematics

### v1

@@ 2.2 SN light-curve simulation and fitting
$S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$ 0.15 3.1 0.05 Table 4 of P21 Table 8 of P21 - -
$S_{\mathrm{C11}}$ 0.15 3.8 0.05 Table 4 of P21 Table 8 of P21 - -
$S_{\mathrm{P23}}$ 0.15 $\mathcal{N}(\mu_{\beta},\sigma_{\beta})$ Table 4 of P21 $\mathcal{N}(\mu_{c},\sigma_{c})$ Eq. 9 Eq. 10
We include in our simulation the reddening effect of the Milky Way dust by using the extinction model of E. L. Fitzpatrick (1999) and the Milky Way reddening map from E. F. Schlafly & D. P. Finkbeiner (2011).
During the simulation process, we apply the magnitude-dependent detection efficiency of the LSST survey as computed in B. O. Sánchez et al. (2022) to each observation. A SN Ia is considered detected if we have at least one observation that passes this detection efficiency. To discard poor-quality SN Ia and to further limit the number of simulated SNe Ia stored, we apply different quality cuts at the end of the simulation stage. We require that the light-curve has at least five observations with a signal-to-noise ratio (SNR) above five and that at least two of them are in different bands amongs the $griz$ bands. In addition, we require at least one observation in the time range of $-20$ to $-5$ days before the SN Ia peak brightness time and one in the time range of 5 days to 40 days after this peak brightness time. To mitigate the effect of dust reddening, we cut all SNe Ia that are observed through the plane of the Milky Way by removing events with a Milky Way dust extinction $E(B-V)$ above 0.25 mag.
@@ 3 Methods
We present here the two approaches used to build the Hubble diagram which we compare in the context of the $f\sigma_{8}$ constraint. The simple framework refers to the approach that was used in C23. The BBC framework refers to the framework originally presented in R. Kessler & D. Scolnic (2017) and improved in P21.
@@ 3.2 The BBC framework
For the $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$ simulations we follow the approach described in P21 as BBC-7D. In this method, the parameters of the SNe Ia from the large simulation are distributed over a 7D grid $\left\{z_{\mathrm{obs}},x_{1},c,\alpha,\beta,\theta,M_{\mathrm{host}}\right\}$. Thus, the bias correction simulations are generated over a $2\times 2$ grid of the parameters $\alpha$ and $\beta$. The extra parameter $\theta$ is a magnitude shift of ±$\pm 0.06$ mag introduced in the bias correction simulation as prescribed in P21. It helps to take into account bias in the recovered $\gamma$ value emerging from correlations between host masses and the SALT stretch and color parameters (M. Smith et al., 2020). Then, bias corrections are computed by averaging in each cell the difference between the estimated SALT parameters and their simulated value,
@@
For the $S_{\mathrm{P23}}$ simulation we use the approach called BBC-BS20 in P21 where the parameters of the grid are reduced to $\left\{z_{\mathrm{obs}},x_{1},c,M_{\mathrm{host}}\right\}$, since $\beta$ has a different value for each SNe and the mass-step is a consequence of the effect of dust on color. The bias correction is directly computed on the difference between the fiducial distance modulus from simulated cosmology and the recovered one
@@
Using the bias correction functions, the nuisance parameters $\alpha$, $\beta$, $\gamma$ and $M_{0,z_{i}}$ along with the intrinsic scatter $\sigma_{\mathrm{int}}$ are determined by the fitting procedure described in J. Marriner et al. (2011), R. Kessler & D. Scolnic (2017) and R. Kessler et al. (2023). In this procedure, the residuals to the fiducial cosmology of the bias correction simulations,
@@
where $\sigma_{\Delta\mu,\mathrm{BBC}}$ is computed for each SN Ia as indicated in Eq. 3 of R. Kessler & D. Scolnic (2017).
In this work we also test the inclusion of the systematic covariance matrix due to the parametrization of the BS21 intrinsic scatter model. To compute this matrix in the BBC framework we generated bias correction simulations to apply to our $S_{\mathrm{P23}}$ data simulation using four variations of our nominal BS21 model parameters. As in V24 we use the parameter values originally found in BS21 and the three random realizations of the parameters drawn from the posterior distributions that were computed in P23. The covariance matrix is computed following the method presented in A. Conley et al. (2010) as
@@
where the $\Delta\mu^{V}_{i}$ are the differences between distance moduli obtained using the nominal bias correction simulations and the ones obtained using the variation $V$ of the BS21 model parameters. The total covariance is then given by
@@ 3.3 Sample cuts
To include a SN Ia in the fit of $f\sigma_{8}$ we need to be able to estimate its host redshift. Hence, after the BBC Hubble diagram fit we require that the SN Ia has been matched to a host using a cut on the dimensionless distance $d_{\mathrm{DLR}}<4$ where $d_{\mathrm{DLR}}$ is defined as the ratio between the angular separation of the SN Ia to its host and the radius of the host in the SN Ia direction (R. R. Gupta et al., 2016; H. Qu et al., 2024). Additionally, in this paper we focus on the $f\sigma_{8}$ fit in the redshift range $z\in[0.02,0.1]$ (a more extensive forecast for different redshift ranges is provided in Rosselli et al. in prep) and therefore cut the SNe Ia which fall outside this range. The lower redshift cut is implemented in order to avoid bias on the velocity estimator, as shown in Appendix A.2 of C23. Averaged over our eight mocks, the final sample size is $\sim 6600$ SNe Ia and is summarized for each intrinsic scatter model simulation in Table 3.
Table 3: Mean number of SNe Ia after cuts over our 8 mocks for each of our intrinsic scatter model simulations in the redshift range $0.10.02<z<$ 0.1.
@@ 3.4 Constraint on $f\sigma_{8}$ using the maximum likelihood method
[…]
@@ 3.4.1 Velocity estimates and covariance matrix
.$\nabla.\mathbf{v}(\mathbf{r},a)=-aH(a)f(a)\theta(\mathbf{r},a),$ (31)
@@
$C^{vv}_{ij}$ $=\langle(\mathbf{v}_{i}.\hat{\mathbf{r}_{i}})((\mathbf{v}_{j}.\hat{\mathbf{r}_{j}})\rangle$ (33)
@@
[…]
@@
These coefficients were fitted on a $N$-body simulation and the accuracy of the resulting power-spectrum is shown to be $>95\%$ for our range of $k$ in J. Bel et al. (2019).
@@ 3.4.2 The redshift space damping function and calibration of the $\sigma_{u}$ parameter
Since positions of galaxies are evaluated using the observed redshift $z_{\mathrm{obs}}$, their distances are affected by velocities through redshift space distortions (N. Kaiser, 1987). To take into account this effect on the velocity covariance we need to introduce a damping function $D_{u}$. An empirical expression for $D_{u}$ has been proposed in J. Koda et al. (2014) and is expressed as
@@
We can see in Fig. 4 that in the case where $f\sigma_{8}$ is fitted along the value of $\sigma_{u}$ seems to converge towards $\sim 21$ Mpc $h^{-1}$. In the case where $f\sigma_{8}$ is fixed we can see that after converging around a value of $\sim 21$ Mpc $h^{-1}$ the $\sigma_{u}$ parameter tends to slightly decrease. A possible interpretation is that with the increase of density, we start to probe more non-linear scales that impact the value of the damping factor. In this work we decide to fix $\sigma_{u}$ to the value of $\sim 21$ Mpc $h^{-1}$. This is comparable to the values ranging from 19 Mpc $h^{-1}$ to 23 Mpc $h^{-1}$ found in Y. Lai et al. (2023) but is slightly higher than the value of $\sim 15$ Mpc $h^{-1}$ found in J. Koda et al. (2014) and C23. We further discuss this systematic in Sect. 4.4.
@@ 3.4.3 The $f\sigma_{8}$ likelihood
We note that in the BBC framework all the $\Theta_{\mathrm{HD}}$ parameters are fixed prior to the $f\sigma_{8}$ fit during the construction of the Hubble diagram. In the case of the simple framework, the $\Theta_{\mathrm{HD}}$ parameters are fit jointly with $\Theta$. We perform the fit by minimizing the negative logarithm of the likelihood of Eq. 41 using the Minuit algorithm (F. James & M. Roos, 1975) as implemented in the Python library iminuit (H. Dembinski et al., 2024).
@@ 4 Results
In this section, we present results obtained after minimizing the negative logarithm of the likelihood from Eq. 41. We focus on estimated PVs and the recovered value of $f\sigma_{8}$. Results obtained for the parameters of the Hubble diagram are presented in the Appendix A.
@@ 4.1 Estimation of PVs for the four scatter models
Fig. 5 presents the residuals of the Hubble diagram and of the estimated velocities for our four simulations stacked over the eight mocks. We can see in the left panel that for the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations, the Hubble diagram residuals have the same scatter of about $\sigma\sim 0.13$ mag for $S_{\mathrm{COH}}$ and $\sigma\sim 0.12$ mag for $S_{\mathrm{G10}}$. In the case of the more chromatic models of simulations $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$ the scatter is again very similar for the two fitting methods with a value of $\sigma\sim 0.13$ mag for $S_{\mathrm{C11}}$ and $\sigma\sim 0.12$ mag for $S_{\mathrm{P23}}$ with a slightly but non-significant better scatter for the BBC fit. These Hubble scatter values result in scatter of estimated PVs in a range between $\sim 1200$ km s-1 and $\sim 1500$ km s-1 for the four different scatter model simulations.
We note in particular that the distribution of the Hubble diagram residuals for the $S_{\mathrm{P23}}$ simulation deviates from the Gaussian one.
@@
This trend is further seen in the pulls of the residuals and velocities in Fig. 6. The pulls of Hubble residuals and velocities are similar for the two fitting methods, except for the $S_{\mathrm{P23}}$ simulation, where we see that the distribution is not Gaussian. In addition to the visual inspection, we compute the skewness $\gamma$ and excess kurtosis $\kappa$ of the distribution ($\gamma=\kappa=0$ for a Gaussian distribution). While the skewness of Hubble residuals is close to zero for the $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$ simulations, for the $S_{\mathrm{P23}}$ simulation we measure $\gamma_{\mathrm{Simp}}\sim 0.54$ for the simple method and $\gamma_{\mathrm{BBC}}=0.34$ for the BBC method. For the kurtosis values we observe that in the Hubble residuals of the $S_{\mathrm{COH}}$ simulation $\kappa_{\mathrm{Simp}}\sim\kappa_{\mathrm{BBC}}\sim 0.09$; in the $S_{\mathrm{G10}}$ simulation $\kappa_{\mathrm{Simp}}\sim\kappa_{\mathrm{BBC}}\sim 0.11$; in the $S_{\mathrm{C11}}$ simulation $\kappa_{\mathrm{Simp}}\sim 0.05$ and $\kappa_{\mathrm{BBC}}\sim 0.04$; and for the $S_{\mathrm{P23}}$ simulation $\kappa_{\mathrm{Std}}\sim 1.6$ and $\kappa_{\mathrm{BBC}}\sim-0.23$. The deviations from zero of the skewness and kurtosis excess for the $S_{\mathrm{P23}}$ are pointing out the presence of non-Gaussianities. These non-Gaussianities in the Hubble diagram distributions are propagated to estimated velocity distributions. This is expected, since the BS21 intrinsic scatter model used in the $S_{\mathrm{P23}}$ simulation has dust-based color dependence that introduces non-Gaussian scatter.
@@
In Fig. 7 we present the velocity residuals as a function of the recovered SALT color parameter $c$. For the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations the two methods of fitting perform similarly. For the two more chromatic models of the $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$ simulations, the simple fit method presents a bias for negative values of $c$. For the $S_{\mathrm{C11}}$ simulation this bias increases linearly below $c\simeq 0$ and goes to $\sim-480$ km s-1 at $c\sim-0.1$. For the $S_{\mathrm{P23}}$ simulation, a similar bias of $\sim-190$ km s-1 is observed for negative $c\sim-0.1$ and an additional bias can be seen for high values of $c$ with an amplitude of $\sim-180$ km s-1 at $c\sim 0.2$. The BBC method avoids this bias through the correction performed in color bins which corrects the average residuals toward a value of zero. Additionally, for the $S_{\mathrm{P23}}$ simulation, the density contours show a correlation between the PV bias and the $c$ parameter; for BBC this correlation seems to be slightly mitigated. However, for both methods, the distribution of the bias on PVs is asymmetric as a consequence of the non-Gaussianity of the residuals.

### v2

@@ 2.2 SN light-curve simulation and fitting
$S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$ 0.15 3.1 0.05 Table 4 of popovic_improved_2021 Table 8 of popovic_improved_2021 - -
$S_{\mathrm{C11}}$ 0.15 3.8 0.05 Table 4 of popovic_improved_2021 Table 8 of popovic_improved_2021 - -
$S_{\mathrm{P23}}$ 0.15 $\mathcal{N}(\mu_{\beta},\sigma_{\beta})$ - Table 4 of popovic_improved_2021 $\mathcal{N}(\mu_{c},\sigma_{c})$ Eq. 9 Eq. 10
We include in our simulation the reddening effect of the Milky Way dust by using the extinction model of fitzpatrick_correcting_1999 and the Milky Way reddening map from schlafly_measuring_2011.
During the simulation process, we apply the magnitude-dependent detection efficiency of the LSST survey as computed in sanchez_snia_2022 to each observation. A SN Ia is considered detected if we have at least one observation that passes this detection efficiency. To discard poor-quality SN Ia and to further limit the number of simulated SNe Ia stored, we apply different quality cuts at the end of the simulation stage. We require that the light-curve has at least five observations with a signal-to-noise ratio (SNR) above five and that at least two of them are in different bands amongs the grizgriz bands. In addition, we require at least one observation in the time range of 20-20 to 5-5 days before the SN Ia peak brightness time and one in the time range of 5 days to 40 days after this peak brightness time. To mitigate the effect of dust reddening, we cut all SNe Ia that are observed through the plane of the Milky Way by removing events with a Milky Way dust extinction E(BV)E(B-V) above 0.25 mag.
@@ 3 Methods
We present here the two approaches used to build the Hubble diagram which we compare in the context of the $f\sigma_{8}$ constraint. The simple framework refers to the approach that was used in carreres_growth-rate_2023. The BBC framework refers to the framework originally presented in kessler_correcting_2017 and improved in popovic_improved_2021.
@@ 3.2 The BBC framework
For the $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$ simulations we follow the approach described in popovic_improved_2021 as BBC-7D. In this method, the parameters of the SNe Ia from the large simulation are distributed over a 7D grid $\left\{z_{\mathrm{obs}},x_{1},c,\alpha,\beta,\theta,M_{\mathrm{host}}\right\}$. Thus, the bias correction simulations are generated over a $2\times 2$ grid of the parameters $\alpha$ and $\beta$. The extra parameter $\theta$ is a magnitude shift of $\pm 0.06$ mag introduced in the bias correction simulation as prescribed in popovic_improved_2021. It helps to take into account bias in the recovered $\gamma$ value emerging from correlations between host masses and the SALT stretch and color parameters (smith_first_2020). Then, bias corrections are computed by averaging in each cell the difference between the estimated SALT parameters and their simulated value,
@@
For the $S_{\mathrm{P23}}$ simulation we use the approach called BBC-BS20 in popovic_improved_2021 where the parameters of the grid are reduced to $\left\{z_{\mathrm{obs}},x_{1},c,M_{\mathrm{host}}\right\}$, since $\beta$ has a different value for each SNe and the mass-step is a consequence of the effect of dust on color. The bias correction is directly computed on the difference between the fiducial distance modulus from simulated cosmology and the recovered one
@@
Using the bias correction functions, the nuisance parameters $\alpha$, $\beta$, $\gamma$ and $M_{0,z_{i}}$ along with the intrinsic scatter $\sigma_{\mathrm{int}}$ are determined by the fitting procedure described in marriner_more_2011, kessler_correcting_2017 and kessler_binning_2023. In this procedure, the residuals to the fiducial cosmology of the bias correction simulations,
@@
where $\sigma_{\Delta\mu,\mathrm{BBC}}$ is computed for each SN Ia as indicated in Eq. 3 of kessler_correcting_2017.
In this work we also test the inclusion of the systematic covariance matrix due to the parametrization of the brout_its_2021 intrinsic scatter model. To compute this matrix in the BBC framework we generated bias correction simulations to apply to our $S_{\mathrm{P23}}$ data simulation using four variations of our nominal brout_its_2021 model parameters. As in vincenzi_dark_2024 we use the parameter values originally found in brout_its_2021 and the three random realizations of the parameters drawn from the posterior distributions that were computed in popovic_pantheon_2023. The covariance matrix is computed following the method presented in conley_supernova_2010 as
@@
where the $\Delta\mu^{V}_{i}$ are the differences between distance moduli obtained using the nominal bias correction simulations and the ones obtained using the variation $V$ of the brout_its_2021 model parameters. The total covariance is then given by
@@ 3.3 Sample cuts
To include a SN Ia in the fit of $f\sigma_{8}$ we need to be able to estimate its host redshift. Hence, after the BBC Hubble diagram fit we require that the SN Ia has been matched to a host using a cut on the dimensionless distance $d_{\mathrm{DLR}}<4$ where $d_{\mathrm{DLR}}$ is defined as the ratio between the angular separation of the SN Ia to its host and the radius of the host in the SN Ia direction (gupta_host_2016; qu_dark_2024). Additionally, in this paper we focus on the $f\sigma_{8}$ fit in the redshift range $z\in[0.02,0.1]$ (a more extensive forecast for different redshift ranges is provided in rosselli_forecast_2025) and therefore cut the SNe Ia which fall outside this range. The lower redshift cut is implemented in order to avoid bias on the velocity estimator, as shown in Appendix A.2 of carreres_growth-rate_2023. Averaged over our eight mocks, the final sample size is $\sim 6600$ SNe Ia and is summarized for each intrinsic scatter model simulation in Table 3.
Table 3: Mean number of SNe Ia after cuts over our 8 mocks for each of our intrinsic scatter model simulations in the redshift range $0.02<z<0.1$.
@@ 3.4 Constraint on $f\sigma_{8}$ using the maximum likelihood method
We fit for $f\sigma_{8}$ using the maximum likelihood method previously used in carreres_growth-rate_2023 and implemented in the Python library flip  [^11] https://github.com/corentinravoux/flip. (ravoux_generalized_2025).
@@ 3.4.1 Velocity estimates and covariance matrix
$\nabla.\mathbf{v}(\mathbf{r},a)=-aH(a)f(a)\theta(\mathbf{r},a),$ (31)
@@
[…]
@@
The linear power spectrum $P_{\theta\theta,\mathrm{lin}}(k)$ is computed using the CLASS library  [^12] https://github.com/lesgourg/class_public. (lesgourgues_cosmic_2011). We take into account smaller scale non-linearities using the formula derived in bel_accurate_2019,
@@
These coefficients were fitted on a $N$-body simulation and the accuracy of the resulting power-spectrum is shown to be $95\%$ for our range of $k$ in bel_accurate_2019.
@@ 3.4.2 The redshift space damping function and calibration of the $\sigma_{u}$ parameter
Since positions of galaxies are evaluated using the observed redshift $z_{\mathrm{obs}}$, their distances are affected by velocities through redshift space distortions (kaiser_clustering_1987). To take into account this effect on the velocity covariance we need to introduce a damping function $D_{u}$. An empirical expression for $D_{u}$ has been proposed in koda_are_2014 and is expressed as
@@
We can see in Fig. 4 that in the case where $f\sigma_{8}$ is fitted along the value of $\sigma_{u}$ seems to converge towards $\sim 21$ Mpc $h^{-1}$. In the case where $f\sigma_{8}$ is fixed we can see that after converging around a value of $\sim 21$ Mpc $h^{-1}$ the $\sigma_{u}$ parameter tends to slightly decrease. A possible interpretation is that with the increase of density, we start to probe more non-linear scales that impact the value of the damping factor. In this work we decide to fix $\sigma_{u}$ to the value of $\sim 21$ Mpc $h^{-1}$. This is comparable to the values ranging from 19 Mpc $h^{-1}$ to 23 Mpc $h^{-1}$ found in lai_using_2023 but is slightly higher than the value of $\sim 15$ Mpc $h^{-1}$ found in koda_are_2014 and carreres_growth-rate_2023. We further discuss this systematic in Sect. 4.4.
@@ 3.4.3 The $f\sigma_{8}$ likelihood
We note that in the BBC framework all the $\Theta_{\mathrm{HD}}$ parameters are fixed prior to the $f\sigma_{8}$ fit during the construction of the Hubble diagram. In the case of the simple framework, the $\Theta_{\mathrm{HD}}$ parameters are fit jointly with $\Theta$. We perform the fit by minimizing the negative logarithm of the likelihood of Eq. 41 using the Minuit algorithm (james_minuit_1975) as implemented in the Python library iminuit (dembinski_scikit-hepiminuit_2024).
@@ 4 Results
In this section, we present results obtained after minimizing the negative logarithm of the likelihood from Eq. 41. We focus on estimated PVs and the recovered value of $f\sigma_{8}$. Results obtained for the parameters of the Hubble diagram are presented in Appendix A.
@@ 4.1 Estimation of PVs for the four scatter models
In the maximum likelihood method, $f\sigma_{8}$ is fitted as the scale of the covariance matrix of the PV distribution, thus it is important to examine this distribution prior to fitting. We represent the distributions of the Hubble diagram residuals and the difference between estimated velocities and the true ones for our four simulations stacked over the eight mocks in Fig. 5. The Hubble diagram scatter has a similar amplitude for the four models, ranging between $\sigma\sim 0.11$ mag for $S_{\mathrm{P23}}$ to $\sigma\sim 0.14$ mag for $S_{\mathrm{C11}}$. The two fitting methods yield the same scatter for the less chromatic models of $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$. For the chromatic models of $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$, the BBC fit yields a slightly but non-significant better scatter than the simple fit. These values of the scatter of the Hubble residuals result in estimated PV scatters in a range between $\sim 1200$ km s-1 and $\sim 1500$ km s-1. We note that for $S_{\mathrm{P23}}$, the distribution of the Hubble diagram residuals deviates from the Gaussian one and seems positively skewed.
@@
This trend is further seen in the pulls of the residuals and velocities in Fig. 6. The pulls of Hubble residuals and velocities are similar for the scatter models of $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$, and $S_{\mathrm{C11}}$ when using both of the two fitting methods. For the $S_{\mathrm{P23}}$ simulation, visual inspection shows that the pull distribution deviates from a Gaussian. To quantify this, we compute the skewness $\gamma$ and excess kurtosis $\kappa$ of the distribution ($\gamma=\kappa=0$ for a Gaussian distribution). Although the skewness of the Hubble residuals distribution is close to zero for $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$, and $S_{\mathrm{C11}}$, for $S_{\mathrm{P23}}$ we measure $\gamma_{\mathrm{Simp}}\sim 0.54$ when using the simple fitting method and $\gamma_{\mathrm{BBC}}=0.34$ when using the BBC fitting method. Similarly, the kurtosis values of the Hubble residuals distribution for $S_{\mathrm{COH}}$, $S_{\mathrm{G10}}$, and $S_{\mathrm{C11}}$ are relatively low, ranging between $\sim$ 0.04–0.12 compared to the value of $\sim 1.62$ found for $S_{\mathrm{P23}}$ when using the simple fitting method. This excess of kurtosis seems to be mitigated when using the BBC fitting method, with a value of $\sim-0.23$. The deviations from zero of the skewness and kurtosis excess in $S_{\mathrm{P23}}$ point to the presence of non-Gaussianities. This is expected, since the brout_its_2021 intrinsic scatter model used in the $S_{\mathrm{P23}}$ simulation has dust-based color dependence that introduces non-Gaussian scatter.
@@
In Fig. 7 we present the velocity residuals as a function of the recovered SALT color parameter $c$. For the simulations $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$, the two fitting methods perform similarly. For the two more chromatic models of $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$, the simple fit method presents a bias when $c<0$. For the simulation $S_{\mathrm{C11}}$, this bias increases linearly below $c\simeq 0$ and reaches a value of $\sim-480$ km s-1 at $c\sim-0.1$. For the simulation $S_{\mathrm{P23}}$, a similar bias of $\sim-190$ km s-1 is observed when $c\sim-0.1$. Furthermore, in $S_{\mathrm{P23}}$, we see a bias for high values of $c$ with an amplitude of $\sim-180$ km s-1 at $c\sim 0.2$. The BBC method avoids this bias through the correction performed in color bins, which corrects the average residuals toward a value of zero. We note that in $S_{\mathrm{P23}}$, the density contours show a correlation between the PV bias and the $c$ parameter; for BBC, this correlation seems to be slightly mitigated. However, for both methods, the distribution of the bias on PVs is asymmetric as a consequence of the non-Gaussianity of the residuals. We further discuss the non-Gaussianity of the HD residuals in Appendix B.